find the time when sum of $2400 increases by 1/5 of every year amounts to $3600.

Additional Details
we need to know the rate of interest as given as per my opinion is the rate is 1/5 or what
Thank you
Joesph

Answer by lwsshak3(11628)

find the time when sum of $2400 increases by 1/5 of every year amounts to $3600.
**
1/5 is the same as 20% interest so we can use the compound interest formula:
A=P(1+i)^n, P=initial investment, i=interest rate per period, A=amount after n periods
..
For given problem:
P=$2400
i=20%
n=to find
A=$3600
..
3600=2400(1+.2)^n
3600/2400=1.2^n
1.5=1.2^n
take log of both sides
log1.5=nlog1.2
n=log1.5/log1.2
n≈2.2 yrs